OpenStudy (anonymous):

the sum of present age of Abe and the age 7 years ago is 31 . Find the present age of Abe. What will be his age after 7 years?

OpenStudy (anonymous):

x+(x-7)=31 2x-7=31 2x=38 x=19 after 7 years he will be.... ??

OpenStudy (anonymous):

19+7= ??

OpenStudy (anonymous):

26 LOL hahaha